jaleah397JK:

Mr. Reddick has 74 total yards of red and blue felt to distribute to students in his art class. Each of his 20 students gets 2.5 yards of blue felt for the project. He also gives each student an equal amount of red felt. How much red felt does each student get? 1.2 yd 3.7 yd 12 yd 24 yd

jaleah397JK:

Please answer

YourMom33:

First off, what do you think it is?

TETSXPREME:

do u have any clue how to do it?

TETSXPREME:

welp o: ik the answer but u needa respond first

jhonyy9:

so like a first step 20*2,5 yards of blue felt = 50 and now subtract this 50 yards from the total yards of 74 and will get the total remained yards of red felt for these 20 students so to get it for each student how much is so just divide the total remained yards of red by 20 hope helped understandably easy @Laylalyssa any idea please ? ty

Laylalyssa:

blue: = 20*2.5 = 50 74-50=24 yards,, red 24/20 = 1.2 every person has to get- 1.2 yards of red felt so then is 1.2 Yards right?